OpenStudy (anonymous):

Which fraction is not equivalent to -5/8? <- Negative is in the middle A. -5/-8 B. -10/16 <- The negative is in the middle

OpenStudy (texaschic101):

what do you mean negative is in the middle ? -5/-8 = 5/8 (not -5/8) -10/16 = - 5/8 (equivalent)

OpenStudy (anonymous):